OpenStudy (anonymous):

what is the slope of -1,4 and -5,2 i need the answer asap

OpenStudy (anonymous):

sloppe of line going through those two points?

OpenStudy (anonymous):

-1-5/ 4-2 = -6/2= -1/3

OpenStudy (anonymous):

if so 1/2

OpenStudy (anonymous):

he's right you add -1 and 5 my bad sorry

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1) = ( 2-4)/(-5+1) = -2/-4 = 1/2

OpenStudy (anonymous):

ok thank you all

OpenStudy (anonymous):

but its a negative 1/2